### v2

# CTR-Driven Ad Text Generation via Online Feedback Preference Optimization

Abstract
Advertising text plays a critical role in determining click-through rates (CTR) in online advertising. Large Language Models (LLMs) offer significant efficiency advantages over manual ad text creation. However, LLM-generated ad texts do not guarantee higher CTR performance compared to human-crafted texts, revealing a gap between generation quality and online performance of ad texts. In this work, we propose a novel ad text generation method which optimizes for CTR through preference optimization from online feedback. Our approach adopts an innovative two-stage framework: (1) diverse ad text sampling via one-shot in-context learning, using retrieval-augmented generation (RAG) to provide exemplars with chain-of-thought (CoT) reasoning; (2) CTR-driven preference optimization from online feedback, which weighs preference pairs according to their CTR gains and confidence levels. Through our method, the resulting model enables end-to-end generation of high-CTR ad texts. Extensive experiments have demonstrated the effectiveness of our method in both offline and online metrics. Notably, we have applied our method on a large-scale online shopping platform and achieved significant CTR improvements, showcasing its strong applicability and effectiveness in advertising systems.
Introduction
Advertising text is crucial in online advertising, appearing in almost every ad. Effective ad texts not only feature well-crafted expression and accurate conveyance of item information (Murakami et al. 2022), but also achieve strong performance in terms of click-through rate (CTR). Large Language Models (LLMs) enable automated ad text generation (ATG) (Mita et al. 2024) from basic item information and simple instructions, offering significantly higher efficiency than manual creation. Some works (Golobokov et al. 2022; Chai et al. 2022; Jin et al. 2023; Kamigaito et al. 2024) focus on offline metrics like diversity or fluency to enhance the quality of generated ad texts. However, these methods do not guarantee higher CTR performance compared to human-crafted texts, revealing a gap between offline metrics optimization and real-world deployment effectiveness.
Figure 1: Challenges in high-CTR ad text generation. (a) Poor diversity and low CTR performance of sampling-based decoding strategy. (b) Noisy online feedback due to uneven page view (PV) distribution across items. (c) Our work propose an innovative two-stage framework: (1) diverse ad text sampling and (2) CTR-driven preference optimization tailored for these challenges.
Recently, studies (Hughes, Chang, and Zhang 2019; Kamigaito et al. 2021; Kanungo et al. 2022; Wei et al. 2022; Murakami et al. 2025) have focused on directly optimizing for online metrics, particularly CTR. CTR has long been regarded as a key metric for evaluating online advertising effectiveness, as clicks serve as the prerequisite for subsequent conversion to other revenue-related metrics. These approaches typically generate multiple ad texts via sampling-based decoding from LLMs for online testing. The generation models are subsequently optimized using online feedback through algorithms such as contrastive learning and reinforcement learning.
However, these approaches often suffer from the low quality of sampling-based decoding and the noise in online feedback, making it difficult to achieve a significant CTR advantage over human-crafted ad texts. We conduct some analytical experiments and reveal the challenges to realize the CTR improvement. Figure 1(a) shows that generating multiple ad texts via sampling-based decoding from LLMs results in poor diversity and low CTR performance. Figure 1(b) illustrates data confidence issues arising from the uneven distribution of item page views (PV) in online feedback. Short-term online testing makes it difficult to accumulate sufficient high-PV data, causing substantial noise in online feedback. These challenges constrain further improvements in CTR optimization of ATG tasks.
In this work, we propose CTOP, a novel CTR-driven ad Text generation method via Online feedback Preference optimization. CTOP employs an innovative two-stage framework consisting of diverse ad text sampling and CTR-driven preference optimization for high-CTR ad text generation, as illustrated in Figure 1(c). In the first stage, we design a diverse ad text sampling strategy to address the low quality of sampling-based decoding. Specifically, for each item, the strategy uses retrieval-augmented generation (RAG) to retrieve multiple chain-of-thought (CoT)-augmented ad text exemplars, each serving as a one-shot prompt to generate a candidate ad text. In the second stage, we conduct CTR-driven preference optimization tailored for noisy online feedback. We select winning candidates that outperform their corresponding human-crafted ad texts in online A/B/n testing and construct preference pairs by treating the former as preferred and the latter as dispreferred. Each pair is assigned a training weight calculated from gain and confidence coefficients based on its CTR improvement and reliability.
Extensive experimental results on offline and online metrics demonstrate the effectiveness of CTOP. Our method enables efficient generation of diverse ad texts and better aligns with real-world user feedback during preference optimization. Notably, we have successfully applied our method to the item title generation task on a large-scale online shopping platform across various advertising display scenarios. Since late 2024, our approach achieved a relative increase of 1.11% in click-through rate (CTR) and 1.02% in revenue per mille (RPM) compared to the human-crafted baseline, showcasing its strong applicability and effectiveness in advertising systems. The main contributions of our work can be summarized as follows:
  • We propose a novel ad text generation method which optimizes for CTR via preference optimization from online feedback, closing the gap between between generation quality and online performance of ad texts.
  • We design an innovative two-stage framework, consisting of (1) diverse ad text sampling via one-shot in-context learning, using RAG to provide exemplars with CoT reasoning; and (2) CTR-driven preference optimization method that dynamically weighs training pairs by gain and confidence coefficients based on online feedback.
  • We demonstrate the practicality and effectiveness of our method through comprehensive offline and online experiments, and show its successful deployment on a large-scale online shopping platform, achieving significant improvements in CTR and RPM (revenue per mille).
Related Works
Advertising Text Generation
Advertising Text Generation (ATG) (Murakami, Hoshino, and Zhang 2023; Mita et al. 2024) aims to automate the creation of high-quality ad texts. Early template-based methods (Bartz, Barr, and Aijaz 2008; Thomaidou et al. 2013) lack diversity and scalability, while extractive approaches (Thomaidou, Leymonis, and Vazirgiannis 2013) ensure content consistency but struggle to generate linguistically varied or user-engaging outputs.
Later works increasingly treat ATG as a sequence-to-sequence task, using pre-trained LLMs to generate fluent, tailored ad texts. These methods explicitly optimize for key dimensions: content diversity (Golobokov et al. 2022; Chai et al. 2022; Jin et al. 2023), fluency (Zhang et al. 2021; Wei et al. 2022; Li et al. 2022), faithfulness (Youngmann et al. 2020; Shao et al. 2021; Golobokov et al. 2022), and relevance to user intent (Kamigaito et al. 2021; Duan et al. 2021). In this work, we directly optimize for the performance metric—click-through rate (CTR)—as the primary objective of ATG.
CTR-Oriented Ad Text Generation
Recently, online performance—particularly click-through rate (CTR)—has emerged as a central objective in ATG research (Wang et al. 2019; Mishra et al. 2020; Kamigaito et al. 2021; Kanungo et al. 2022; Wei et al. 2022; Murakami et al. 2025), as it directly reflects user engagement and advertising effectiveness. Typically, these studies first generate diverse ad texts via sampling-based decoding for online A/B testing. Then, different algorithms are leveraged to learn from online feedback, such as reinforcement learning (Hughes, Chang, and Zhang 2019; Kamigaito et al. 2021; Kanungo et al. 2022), contrastive learning (Wei et al. 2022), semi-supervised learning (Shao et al. 2021), and training-free in-context learning (Murakami et al. 2025).
However, these methods suffer from the low-quality sampling-based decoding and the noisy online feedback, leading to insufficient alignment with user preference. In our work, we design diverse ad text sampling and CTR-driven preference optimization tailored for these issues.
Reinforcement Learning in Advertising Creativity
With the success of Reinforcement Learning from Human Feedback (RLHF) (Christiano et al. 2017; Ziegler et al. 2019; Ouyang et al. 2022; Ji et al. 2023; Wang et al. 2024), RL has been widely used in ad creativity tasks like text and image generation. Early works (Hughes, Chang, and Zhang 2019; Kamigaito et al. 2021; Kanungo et al. 2022) adopt self-critical RL (Rennie et al. 2017) to optimize predicted CTR (pCTR) or proxy metrics like ad quality scores. Recent works in ad image generation (Yang et al. 2024b; Chen et al. 2025) train CTR reward models on user preferences and use them to guide PPO (Schulman et al. 2017) or label preference data for DPO (Rafailov et al. 2023). In our work, we propose CTR-driven preference optimization based on the DPO algorithm to optimize ATG for CTR improvement.
Task Formulation and Challenges
CTR-driven Ad Text Generation
We formulate the Ad Text Generation (ATG) task with click-through rate (CTR)—the core online metric—as the optimization objective. Given the item information $x$, the goal is to learn a parameterized generator $G_{\theta}$ that maximizes the expected CTR of generated ad texts $y$:
$\theta^{*}=\arg\max_{\theta}\mathbb{E}_{x\sim\mathcal{X},\,y\sim G_{\theta}(\cdot|x)}\left[\mathrm{ctr}(y)\right],$ (1)
where $\mathrm{ctr}(y)=\frac{\mathrm{click}(y)}{\mathrm{pv}(y)}$ is computed from online clicks and ad page views ($PV$).
However, evaluating whether a generated ad text achieves “high CTR” is challenging without a meaningful baseline. In practice, each item has a human-crafted ad text $y_{\text{hu}}$ typically written by merchants or the advertising platforms based on domain expertise or empirical knowledge. These human-crafted ad texts generally achieve a certain level of quality and CTR, which serves as a strong and interpretable baseline. Therefore, any generated ad text $y$ that outperforms $y_{\text{hu}}$ in online testing can be considered achieving high CTR. To this end, we reformulate the objective as maximizing the CTR improvement over human-crafted ad texts:
$\theta^{*}=\arg\max_{\theta}\mathbb{E}_{x\sim\mathcal{X},\,y\sim G_{\theta}(\cdot|x)}\left[\mathrm{ctr}(y)-\mathrm{ctr}(y_{\text{hu}})\right].$ (2)
This relative formulation is more tractable to optimize and better business-aligned than the absolute form in Equation 1. It calls for the effective generation of candidate ad texts with high-CTR potential to outperform human-crafted baselines during online testing.
Figure 2: Item and candidate win rates over human-crafted ad texts with different numbers of candidates per item ($K$).
[…]
$n=1$ $n=2$ $n=2$ $n=3$
5 0.540 0.376 0.422 0.712
10 0.680 0.497 0.342 0.650
Table 1: Diversity metrics of ad texts generated by sampling-based decoding ($K=5$ and $K=10$ candidates per item).
Challenge I: Low Quality of Sampling-based Decoding
We select the item title generation task on a large online shopping platform to explore generating high-CTR candidate ad texts. Given the flexibility of linguistic expression, introducing diversity into candidate ad texts is a natural strategy to generate more high-CTR candidates.
[…]
We define the item win rate as the proportion of items for which at least one candidate title achieves a higher CTR than the human-crafted. As shown in Figure 2, the item win rate increases with $K$, indicating that introducing diversity contributes to generating more high-CTR ad texts.
[…]
Sampling-based decoding restricts generation to high-probability tokens but still inherits model biases and risks output degradation, especially at high temperatures. The poor diversity and inferior CTR performance reveal the low quality of sampling-based decoding.
Figure 3: Distribution of item page views ($PV$) after two weeks of online testing.
Figure 4: Overall framework of our CTR-driven ad text generation method, CTOP. (a) diverse ad text sampling via one-shot in-context learning, using RAG to provide exemplars with CoT reasoning. (b) Online A/B/n testing and a parallel control AA group with identical settings for confidence measurement. (c) CTR-driven preference optimization weighing preference pairs according to their CTR gain and confidence levels.
$PV$ Level Low Medium High
Pair Accuracy (%) 52.3 54.9 57.6
Table 2: Pair accuracies of DPO models trained with preference data of different item page view ($PV$) levels.
Challenge II: Optimization under Noisy Feedback
Optimizing ad text generation for high CTR aligns naturally with reinforcement learning, where the goal is to maximize user engagement in real-world environments. In our setting, online A/B/n testing offers an off-policy dataset comprising both LLM-generated and human-crafted texts. Moreover, CTR feedback is delayed, noisy, and requires significant online accumulation, making it challenging to train an explicit reward model for CTR prediction or preference estimation. These factors motivate us to adopt off-policy RL methods with implicit reward modeling (Rafailov et al. 2023; Pal et al. 2024; Wu et al. 2024; Azar et al. 2024; Kim et al. 2024; Ethayarajh et al. 2024; Liu et al. 2025).
Here we adopt DPO (Rafailov et al. 2023) due to its simplicity and efficiency. DPO operates on preference tuples $(x,y_{w},y_{l})$, where $y_{w}$ and $y_{l}$ are the preferred and dispreferred responses with input context $x$. Given a policy model $\pi_{\theta}$ and a reference model $\pi_{\mathrm{ref}}$, the DPO objective is:
$\mathcal{L}_{\text{DPO}}=-\log\sigma\left(\beta\left[\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right]\right)$ (3)
where $\beta$ denotes the preference strength, and $\sigma$ is the sigmoid function.
We construct preference pairs with each winning candidate as $y_{w}$ and its human-crafted ad text as $y_{l}$. A Qwen-2.5-7B model is supervised fine-tuned on diverse ad texts as $\pi_{\mathrm{ref}}$. With a dataset of 100,000 items from online A/B/n testing, we train a DPO-optimized model. However, A/B testing against human-crafted ad texts yields only a 41.2% win rate.
We attribute this to the low confidence of preference pairs, especially for items with limited page views ($PV$). To investigate, we conduct an offline study by splitting the preference data into three PV-level subsets: low ($PV\leq 100$), medium ($100<PV\leq 500$), and high ($PV>500$). A separate high-PV test is constructed to evaluate pairwise accuracy:
$\text{Acc}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[\pi_{\theta}(y_{w}^{(i)}\mid x^{(i)})>\pi_{\theta}(y_{l}^{(i)}\mid x^{(i)})\right]$ (4)
As shown in Table 2, preference data with higher $PV$ yields better DPO performance due to its higher confidence. However, Figure 3 presents the uneven item $PV$ distribution after two weeks of online testing, with only a small fraction reaching high $PV$. Training data from noisy online feedback limits the effectiveness of preference optimization.
Method
In this section, we introduce CTOP, a novel CTR-driven ad Text generation method via Online feedback Preference optimization. To ensure meaningful improvements in online performance, we formulate the task objective as maximizing the CTR improvement over human-crafted ad texts, defined in Equation 2. As illustrated in Figure 4, our method employs an innovative two-stage framework of diverse ad text sampling and CTR-driven preference optimization tailored for the two challenges. To address the low-quality issue of sampling-based decoding, we design a diverse ad text sampling strategy in the first stage. To better utilize noisy online feedback, we perform CTR-driven preference optimization in the second stage. The following subsections elaborate on each stage of our proposed method.
Diverse Ad Text Sampling
As depicted in Challenge I, candidate ad text generation via sampling-based decoding causes poor diversity and inferior CTR performance. This motivates the design of a strategy that achieves both high quality and true diversity.
[…]
However, direct imitation of retrieved ad text often fails to capture their stylistic essence. We prompt the model to first analyze the retrieved exemplar and extract imitation tips before ad text generation. This chain-of-thought (CoT) reasoning process yields improvements in the quality of the generated ad texts. To address the inefficiency of online CoT during inference, we pre-generate offline CoTs for all exemplars in the repository. By retrieving exemplars with their CoTs, the in-context learning maintains quality while speeding up ad text generation.
Owing to the inherent stylistic diversity among high-quality exemplars in the repository, our strategy is able to maintain both quality and diversity even as the number of candidates per item $K$ increases. Moreover, this diversity is controllable: our strategy can generate new types of ad texts simply by collecting additional exemplars without extra training. Compared to methods that demand manual prompt engineering for each new style, our approach offers a more scalable and efficient solution with minimal overhead.
CTR-Driven Preference Optimization
Following our sampling strategy, we deploy the generated candidates and the human-crafted ad texts for short-term online A/B/n testing, as shown in Figure 4(b). A candidate that achieves higher CTR than its corresponding human-crafted ad text is defined as a winning candidate. We construct preference pairs with each winning candidate as $y_{w}$ and its human-crafted ad text as $y_{l}$. However, our analysis in Challenge II reveals that noisy online feedback negatively impacts the quality of preference optimization. To address this issue, we propose a CTR-driven preference optimization approach illustrated in Figure 4(c), where each pair is assigned a dynamic weight calculated from gain and confidence coefficients.
The gain coefficient measures the improvement of the winning candidate’s CTR over its human-crafted ad text. We first compute the absolute CTR difference for each preference pair within the same item. Let $\Delta_{i}^{(k)}=\mathrm{ctr}\big{(}y_{w}^{(k)}\big{)}-\mathrm{ctr}\big{(}y_{l}^{(k)}\big{)}$ denote the CTR difference within the $k$-th pair of item $i$. To account for varying baseline CTRs across items and ensure numerical stability, we normalize $\Delta_{i}^{(k)}$ by the maximum CTR difference within the same item:
$\mathrm{gain}_{i}^{(k)}=\frac{\Delta_{i}^{(k)}}{\max_{k^{\prime}}\Delta_{i}^{(k^{\prime})}}\times m$ (5)
where $m$ is a scaling factor that amplifies small differences for better distinguishability in preference optimization.
The confidence coefficient quantifies the reliability of each pair, based on the relative order of the pair between the experimental group and the parallel control group with identical settings. As shown in Figure 4(b), the parallel control group mirrors the experimental group setup exactly, including ad content and traffic allocation. In online advertising, such a parallel control group is commonly referred to as the AA group, serving to evaluate the stability of ad performance. For each preference pair $(y_{w},y_{l})$, we compare its relative order in both the experimental group and the AA group for confidence calculation:
$\mathrm{confidence}=\begin{cases}1,&y_{w}\succ_{\text{exp}}y_{l}\cap y_{w}\succ_{\text{AA}}y_{l},\\ 0.5,&\text{otherwise}.\end{cases}$ (6)
This binary confidence score ensures that stable and reproducible preferences contribute fully to model training, while mitigating the impact of unreliable pairs.
For each preference pair $k$ of item $i$, we assign a dynamic weight $w_{i}^{(k)}=\mathrm{gain}_{i}^{(k)}\cdot\mathrm{confidence}_{i}^{(k)}$. Building upon DPO (Rafailov et al. 2023) algorithm, our final CTR-driven preference optimization objective is formulated as follows:
$\mathcal{L}_{\text{CTRPO}}=-\sum_{i=1}^{N}\sum_{k=1}^{M_{i}}w_{i}^{(k)}\cdot\mathcal{L}_{\text{DPO}}$ (7)
where $N$ is the total number of items, $M_{i}$ is the number of preference pairs constructed for item $i$, and $\mathcal{L}_{\text{DPO}}$ denotes the DPO loss as formulated in Equation 3.
Our approach enables prioritizing preference signals of both substantial CTR improvement and high reliability, leading to more robust and effective ad text generation.
Experiments
In this section, we primarily evaluate our method through online experiments, supplemented by offline studies. We first compare our proposed method with other baselines via short-term online A/B testing with human-crafted ad texts. Next, we conduct experiments to validate the impact of the two stages in the framework, respectively. Finally, we present the long-term online business gains by deploying our method at scale on a large-scale online shopping platform.
Experiment Setup
Implementation Details
We conduct online experiments on a large-scale e-shopping platform, with the task of generating item titles across display advertising scenarios. For diverse ad text sampling, we first build a high-quality item title repository with offline CoTs. Then we retrieve the top-5 semantically relevant exemplars per item and use each as a one-shot prompt of style imitation. A pre-trained Qwen-2.5-7B is used to generate 5 candidate titles for the current item through in-context learning. We select 200,000 items across different categories for diverse generation, resulting in a total of 1 million ad titles. These candidates, together with the human-crafted item titles, are deployed for a two-week online A/B/n testing using 5% of platform traffic as the experimental group. A parallel AA group is simultaneously deployed with identical content and traffic allocation. Due to business sensitivity concerns, we are unable to publicly release the online testing data used in this study.
For CTR-driven preference optimization, we first construct a preference dataset of 500K pairs by pairing each winning candidate and its corresponding human-crafted title from the online A/B/n testing. Each pair is assigned a dynamic weight based on its gain coefficient (Equation 5, with scaling factor $m=2.5$) and confidence coefficient (Equation 6). We use a Qwen-2.5-7B model supervised fine-tuned on diverse titles in the A/B/n testing as the reference model $\pi_{\text{ref}}$, which also initializes the policy model $\pi_{\theta}$. We optimize the objective in Equation 7 for one training epoch with a learning rate of $2\times 10^{-5}$. All training processes are conducted on machines equipped with 8 NVIDIA A100 GPUs.
Baselines
For fair comparison, we evaluate our method against the following baselines with the same online data by our diverse ad text sampling strategy. All baseline models are implemented with Qwen-2.5-7B.
  • Prompt-only: Directly generating titles from item information using simple prompts without extra model tuning.
  • Top-1 sampling: Our diverse ad text sampling with sampled number $K=1$.
  • Top-1 SFT: Supervised fine-tuning on the highest-CTR winning candidate of each item from the A/B/n testing.
[…]
Evaluation Metrics
We select 1 million items for evaluation, ensuring that none of them overlap with those in the training set. For each method, a single ad title is generated for each item then deployed alongside its corresponding human-crafted title for a two-week A/B testing on the online shopping platform, using 5% of total traffic. The following metrics are calculated from the online feedback:
  • Win Rate: The proportion of items where the generated title achieves a higher CTR than the human-crafted title.
  • Overall Relative CTR Improvement (CTR $\uparrow$): The relative improvement in average CTR of the generated titles over the human-crafted titles, aggregated across all items.
Method Win Rate (%) CTR $\uparrow$ (%)
training-free
Prompt-only 32.1 -9.83
Top-1 sampling 48.7 -0.6
preference alignment
Top-1 SFT 52.3 +0.92
DPO 56.0 +3.48
CTOP (Ours) 60.2 +4.76
Table 3: Comparison of online A/B testing results across different methods. Our method outperforms all baselines.
Figure 5: Win rate (over human-crafted ad texts) comparison between sampling-based decoding and our diverse ad text sampling with different numbers of candidates ($K$).
Figure 6: Comparison cases between human-crafted item titles and our generated titles on the online shopping platform.
Main Results
Table 3 shows the online A/B testing results. Our method achieves the best win rate (60.2%) and relative CTR improvement (+4.76%), outperforming all baselines and demonstrating a significant positive impact on overall CTR.
For training-free baselines (Prompt-only and Top-1 sampling), both underperform the human-crafted ad texts. Prompt-only yields a low win rate (32.1%) and negative CTR improvement (-9.83%), revealing the gap between model generation quality and online deployment effectiveness. Top-1 sampling improves win rate through only the diverse ad sampling strategy, but still results in a slight CTR reduction (-0.6%). This suggests that direct imitation does not ensure suitability for the target item, underscoring the need for preference optimization from online feedback.
In contrast, preference alignment methods leverage online feedback for model optimization, all achieving higher win rates ($>$50%) over the human-crafted baseline. Among these methods, DPO outperforms Top-1 SFT, demonstrating the advantage of reinforcement learning to align with user engagement. Our method CTOP extends DPO by introducing dynamic weights of preferences based on CTR gain and confidence, achieving superior performance. Through prioritizing high-confidence, high-gain pairs, the model learns from the most meaningful feedback, leading to more robust and effective ad text generation.
#Cand.
($K$)
[…]
$n=1$ $n=2$ $n=2$ $n=3$
5
Sampling-based
decoding
0.540 0.376 0.422 0.712
Ours 0.454 0.277 0.648 0.930
10
Sampling-based
decoding
0.680 0.497 0.342 0.650
Ours 0.599 0.389 0.555 0.890
Table 4: Comparison of diversity metrics between sampled-based decoding and our diverse ad text sampling ($K=5$ and $K=10$ candidates per item).
Analysis on Diverse Ad Text Sampling
The comparison between Prompt-only and Top-1 sampling in the main results (Table 3) has demonstrated the effectiveness of our diverse ad text sampling strategy. To provide a further analysis, we comprehensively compare our method with sampling-based decoding ($temperature=1.5,top\_k=100$) across different candidate numbers $K$.
As shown in Figure 5, our method consistently outperforms sampling-based decoding on both item and candidate win rates across all $K$. Furthermore, while the candidate win rate of sampling-based decoding drops with larger $K$ (from 41.0% to 25.9%), our method remains stable at around 50%, indicating strong quality of the individual candidate.
[…]
Confidence Gain Win Rate (%) CTR $\uparrow$ (%)
56.0 +3.48
58.5 +3.79
60.2 +4.76
Table 5: Ablation results of gain/confidence coefficients in CTR-driven preference optimization.
Ablation on CTR-driven Preference Optimization
We conduct an ablation to evaluate the impact of gain and confidence coefficients in our CTR-driven preference optimization. As shown in Table 5, removing both coefficients reduces the method to standard DPO, which yields a 56.0% win rate and +3.48% relative CTR improvement. Adding only the confidence coefficient improves the win rate to 58.5% and CTR to +3.79%, indicating better stability by prioritizing reliable preference pairs. Our full method, which incorporates both gain and confidence, further boosts the win rate to 60.2% and relative CTR improvement to +4.76%. These results demonstrate that jointly considering both factors produces the best overall performance: gain emphasizes the impact of preference pairs on CTR, while confidence mitigates noise from unreliable data.
Online Results
We deployed our proposed CTOP for item title generation on a large-scale e-shopping platform across various ad scenarios. In initial A/B/n testing, the candidate group showed a 0.46% CTR decrease versus the human-crafted baseline, indicating the necessity of preference alignment. After CTR-driven optimization, the generated titles achieved a relative improvement of +1.11% in CTR and +1.02% in RPM (revenue per mille) over the human-crafted baseline under full platform traffic, since late 2024. The significant performance improvement demonstrates the strong applicability and effectiveness of CTOP in real-world advertising systems.
Case Study
Figure 6 provides comparison cases between human-crafted item titles and our generated titles on the shopping platform. The human-crafted titles are often lengthy, keyword-stuffed, lacking readability or user appeal. In contrast, our generated titles are concise and stylistically diverse—ranging from benefit-driven to persona-targeted—using engaging language that is well-aligned with both item features and user attention, resulting in better CTR performance.
Conclusion
In this work, we propose CTOP, a novel CTR-driven ad text generation method via online feedback preference optimization. CTOP employs a two-stage framework, comprising a diverse ad text sampling and a CTR-driven preference optimization approach, tailored for the low-quality sampling-based decoding and noisy online feedback, respectively. Extensive experiments have demonstrated the effectiveness of our method in both offline and online metrics, significantly improving CTR in real-world advertising scenarios. This study contributes an effective solution to the industry and inspires further research in e-commerce advertising.